Five times a certain number, X, exceeds 8 by three times the sum of the number and 2. Find the number.
Answer by ankor@dixie-net.com(22740)

Five times a certain number, X, exceeds 8 by three times the sum of the number and 2.
Find the number.
5x = 8 + 3(x+2)
:
5x = 8 + 3x + 6
:
5x - 3x = 8 + 6
:
2x = 14
;
x = 7 is the number
:
Check in the original statement
5(7) = 8 + 3(7+2)
35 = 8 + 27